# AI Failure Loops in Devalued Work: The Confluence of Overconfidence in AI & Underconfidence in Worker Expertise

### v1

@@ 1 Introduction
Workplace AI systems are often touted by developers and deployers as tools to augment workers’ abilities and improve their work experience (Brynjolfsson et al., 2025). In practice, however, these systems have frequently fallen short of meaningfully supporting workers (Acemoglu et al., 2023). In social services, AI-based decision support tools have been introduced for high-stakes decisions, only to be withdrawn following backlash from frontline workers (Samant et al., 2021). In healthcare, clinicians ignore AI tools designed to assist in clinical decision-making (Elwyn et al., 2013; Lebovitz et al., 2022). In education, AI tutoring systems have added to teachers workloads and compromised their pedagogical autonomy—despite promises to reduce labor and enhance teaching (Holstein et al., 2017; Utterberg Modén et al., 2021; Holmes, 2023). Across these sectors, AI systems fail to complement the situated, social, and tacit knowledge that underpins workers’ expertise (Lebovitz et al., 2021).
Prior research in human-computer interaction (HCI) and science and technology studies (STS) offers several explanations for these failures. Scholars have pointed to issues of poor contextual fit, where AI systems conflict with existing organizational structures and work practices (Suchman, 1987; Yang et al., 2019). Others have highlighted flawed assumptions about the nature of work, misaligned values, or inadequate governance, emphasizing how these design shortcomings intersect with broader labor and power dynamics  (e.g., Suchman, 1993; Star, 1999; Kawakami et al., 2024a). However, one crucial dimension remains underexplored: occupational devaluation, the systemic undermining of worker expertise, particularly within professions predominantly occupied by women and people of color (Flores-Robles and Gantman, 2023). In these professions, expertise is often not fully understood or explicitly recognized by others (Huising et al., 2025), and devalued both in status and pay even when accounting for education and experience levels (Flores-Robles and Gantman, 2023).
@@
We argue that workers engaged in devalued labor are especially vulnerable to flawed AI deployments, which often rest on reductive understandings of their work. These systems not only fail to reflect the complexity of frontline labor but can also further obscure and erode the visibility of workers’ expertise, creating a negative feedback loop that deepens occupational devaluation. Drawing together past scholarship on AI deployments across three feminized labor contexts (social work, home healthcare, K-12 teaching), we articulate why and how this dynamic emerges.
To do so, we introduce the concept of AI Failure Loops: recurring patterns of failure in the design, evaluation, and deployment of workplace AI systems, in which the devaluation of worker expertise leads to ineffective AI deployments—and those failed deployments, in turn, further erode recognition of workers’ skills. These loops are sustained across the AI development lifecycle, fueled by persistent over-estimation of what AI can do and under-recognition of the complexity, nuance, and value of worker expertise.
With recent waves of both AI hype and actual increases in AI capabilities, developers are targeting an expanding range of tasks associated with feminized occupations. Drawing lessons from past AI deployments in these contexts, we can expect this will further accelerate the gross under-estimations of worker expertise and over-estimations of AI capabilities that has shaped these occupations in the past three decades. This paper responds to existing calls to redirect this trajectory toward pro-worker futures (e.g., Acemoglu et al., 2023), where worker expertise is celebrated within AI practice and society more broadly, and workplace AI systems are designed to truly complement worker capabilities. By centering occupational devaluation in our analysis, we extend these calls, highlighting the need for AI systems that not only avoid harm but actively contribute to the dignity and visibility of labor.
@@ 3 Devaluation of Feminized Labor
Notions of what counts as “work” and, relatedly, expertise are socially situated and have evolved across time (Suchman, 1995; Star and Strauss, 1999). These notions are used to justify the societal and economic (de)valuation of human labor. Feminized labor—which often involves care-oriented labor like teaching, home healthcare, or childcare—provides a particularly striking example. Many forms of feminized labor were historically not viewed as “work,” but as acts of love (Flores-Robles and Gantman, 2023). Even as feminized labor became formalized into paid occupations through social movements, workers struggled to gain legitimacy. Sociologists have theorized several explanations: Employers may justify low pay by conceptualizing feminized labor as “acts of love” provided out of “goodwill” (the “prisoners of love” framework (England, 2005)). The association of feminized labor with women may also lead to its devaluation precisely because women themselves are devalued in society (the devaluation theory) (England, 2005).
Even when accounting for factors like workers’ education level, occupations with a higher percentage of women pay less and have lower social status (Flores-Robles and Gantman, 2023). The most female-dominated occupations in the United States still tend to involve care-based work, including home care (85%), social work (85%), and teaching at the primary and secondary levels (77%) (Council on Social Work Education, 2023; Connacher, 2023; Taie and Lewis, 2022). Workers in feminized occupations tend to have low status in their workplace, with little say in organizational decisions impacting their day-to-day work—challenges which are exacerbated by a “glass ceiling” on career advancements (Cotter et al., 2001). These occupations often also face systemic resource shortages and high rates of worker burnout, all amid growing labor demands (England, 2005).
@@ 4.1 Team Experience and Case Selection
We examine case studies of AI deployments across social work, home healthcare, and K-12 teaching, selected because our team has direct experiences collaborating with workers in each occupation. This experience—collectively including over ten years of work observing worker practices, designing interventions with workers, and speaking with leadership and managers to facilitate these collaborations—-allowed us to ground our analysis of relevant literature in our own empirical material (e.g., field notes). For example, our social work case study was informed by our prior findings from field research conducted across five U.S. public sector agencies (Kawakami et al., 2022a, 2024a). This includes experiences observing workers’ use of AI-based risk assessment tools (Kawakami et al., 2022a), speaking with agency leadership and in-house AI developers in social work agencies (Kawakami et al., 2024a, b), and designing interventions with social workers (Kawakami et al., 2022b). Our healthcare case study was informed by our experiences conducting interviews and design activities with home health aides, aide coordinators, and nurses to understand their perspectives around workplace AI (e.g., Bartle et al., 2022). Our K-12 teaching case study was informed by our prior field research studying K-12 teachers’ adoption and use of AI-based tutoring systems across hundreds of US-based schools and classrooms (e.g., Holstein et al., 2017, 2018, 2019).
@@ 4.2 Literature Review and Analysis
As we conducted a comparative case analysis (Schwandt and Gates, 2018) of AI deployments across these occupations, we began to notice common trends, which we explored further through a focused review of academic and grey literature. We gathered and analyzed literature from computing (e.g., HCI, robotics, AIED) and domain-specific (e.g., healthcare, social work, and education) research venues. We searched literature on field deployments of AI systems in workplace contexts (e.g., Jørgensen and Nissen, 2022), literature reviews on technologies deployed in specific workplaces (e.g., Kuo et al., 2022), and syntheses of trends on the impacts of AI on workers (e.g., Holmes, 2023). Our analysis was not limited to research on challenges from AI deployments. Through our initial comparative case analysis of findings across the authors, and this subsequent review of the literature, we iteratively developed the AI Failure Loops concept.
@@ 6 Case Studies
Considered part of domestic life for most families in early 19th century, while professional home care was reserved for wealthy families. By late 19th century, women began organizing efforts to send nurses to homes of poorer families (Buhler-Wilkerson, 2003). The Fair Labor Standards Act of 1938 outlaws exploitative practices and instituted an 8-hour workday but home health aides are explicitly excluded (Migda, 2015). Home health aides continue to be predominantly women and disproportionately Black or Hispanic (Connacher, 2023). Continued worker devaluation: Majority of workers lack benefits like paid time off, sick leave, or employer-sponsored health insurance (Banerjee et al., 2021); majority of workers live below poverty line (Connacher, 2023); workers often not seen as part of healthcare team, despite being integral to healthcare system (Osterman, 2018).
@@
Originally conceived as male-dominated occupation that men partook in during farming off-seasons or prior to moving into more well-established professions. Women begin to enter the profession in the late 1880s (Rury, 1991). Profession is predominantly occupied by women, except among administrators (Smith, 2022). Profession becomes increasingly formalized with state licensing and university education program requirements (tea, 2018). K-12 teachers continue to be predominantly women, except in leadership positions (e.g., principals, superintendents) which continue to be male-dominated (Kocon, 2017). Continued worker devaluation: Paid relatively little compared to other fields requiring similar levels of education (Flores-Robles and Gantman, 2023); workers report high levels of stress and burnout (Steiner and Woo, 2021).
@@ 6.1 AI-Based Risk Assessment Tools Have Amplified Misconceptions of Social Workers’ Expertise
AI Failures in Social Work
@@
Context
Governments and private investors have thrown billions of dollars into healthcare AI innovations—and this number is only increasing each year (Esposito et al., 2016; Yang et al., 2019; Yildirim et al., 2023). For example, to help address the immense care work shortage, several countries–across Europe, North America, and East Asia–have heavily invested in exploring the potential of care robots (Esposito et al., 2016). In fact, care robots for home healthcare applications are among the main investments in robotics applications (Esposito et al., 2016). However, a growing body of work has recognized that, if not designed carefully, these systems may disempower workers and amplify challenges they face already (Fox et al., 2019; Tseng et al., 2020; Persson et al., 2022). Today, there exist few, if any, technical innovations that actually support home health aides (Kuo et al., 2022)—workers, the majority of whom are women and people of color, who support elderly people and people with disabilities in everyday personal care (e.g., bathing, washing, monitoring health challenges) and household responsibilities (e.g., cooking, cleaning, shopping). This has occurred amid increasing acknowledgement that home health aides are the cornerstone to systems of care, that demands for their labor are increasing, and that institutional support for their wellbeing and labor has remained stagnant (Connacher, 2023).
AI Failures in Healthcare
@@
Even as decision-makers in healthcare (e.g., doctors, hospitals, state regulators, insurance companies) verbally acknowledge home care aides, they “typically do not think that home care aides can be real members of a care team”  (Osterman, 2019). Osterman describes this happens, in part, because home care aides’ skills are overlooked, instead misunderstood as that of “unskilled companions [and] glorified babysitters” (Osterman, 2019).
@@
Researchers and policymakers have long looked for ways to sustain the growing population of elderly people—much of which is currently supported through home health aides (Bedaf et al., 2015; Hendry and Wu, 2004). While there has been limited research on technologies to support home health aides, there is a comparatively large number of papers exploring ways to automate home-based tasks needed by elderly people, including tasks home health aides usually complete (Esposito et al., 2016). Many of these technical deployments have come in the form of care robots. However, despite the massive amounts of funding dedicated towards advancing the design of care robots, actual deployments of care robots have often been deemed failures (e.g., Persson et al., 2022; Sharkey and Wood, 2014). Most of this literature has discussed failures with respect to meeting elderly people’s care needs. In this section, we discuss how care robots have also failed to bring value to home health aides themselves.
While care robots may have been intended to help alleviate the increasing demand for home care workers, several studies have observed impacts in the opposite direction. For example, in a scoping review of 27 papers on the deployment of care robots, researchers concluded the effectiveness of care robots is “mixed” at best. It further observed how robots can instead increase caregivers’ workload: care robots “seldom work as a shortcut to increased efficiency or time effectiveness” as workers must undergo the labor of maintaining and operating the robot (Persson et al., 2022). Another study found care workers experienced emotional demands, grappled with new ethical challenges, and felt decreased job satisfaction when care robots were deployed in their workplace (Wright, 2018, 2019).
@@
Home health aides are rarely consulted during the design process, even for technologies directly impacting them (Kuo et al., 2022). This creates a stark contrast with the development of technologies that support doctors. In this section, we briefly discuss how having organizational power as a “high status” worker can create incentives to avoid the AI Failure Loop. In doing so, we shed light onto the power and labor dynamics at play in low status occupations, that lead to deviations in how they are impacted by AI deployments.
@@ 6.3 Persistent Visions of Automation in Education Have Further Devalued Teachers’ Expertise
There is a long history of AI development in education guided by the ultimate vision of automating away the role of human teachers (Watters, 2015; Selwyn, 2016). Over the past forty years, research and practice in AIED has largely centered students’ needs, often without explicit consideration for teachers’ needs and constraints in the classroom (Baker, 2016; Holmes, 2023; Holstein et al., 2017; Roll and Wylie, 2016). Holmes refers to these deployments as “student-focused AIED”—AI technologies aiming to automate one or more of teachers’ tasks for the stated goal of helping students learn, for example, including tutoring systems, virtual writing assistants, automatic writing evaluation, and chatbots. By contrast, Holmes finds that “teacher-focused AIED” has remained largely speculative, confined to academic research prototypes rather than commercial systems (Holmes, 2023). Beyond the research community, commercialization practices help explain the prevalence of student-focused AIED technologies at the expense of teacher-focused AIED. Companies sell AIED technologies to schools and districts, and past research has found that teachers may be given little say in whether or not to adopt a given technology (Holstein et al., 2017; Utterberg Modén et al., 2021).
By reducing K-12 teachers to mere service providers responsible for operating technology and maintaining the classroom (Holmes, 2023), many commercialized AIED deployments require teachers to “accommodate” to the technology. Indeed, prior work examining AI deployments observed how teachers often struggle to make AIED technologies work in their classroom, without having to adjust their practices in unreasonable ways or take on new responsibilities along the way (Enyedy, 2014; Utterberg Modén et al., 2021). Observing math teachers using an AI tutoring systems in their classrooms, Holstein et al. and Utterberg Moden et al. found that teachers were forced to grapple with numerous contradictions designed into the AIED system, eventually causing them to abandon the system. For instance, while the AI tutoring system aimed to help students work at their own pace, this ran counter to pedagogical practices around collaborative learning and discussion: Teachers had a responsibility to maintain a cohesive classroom environment, where students were roughly synchronized in the content they were learning. (Holstein et al., 2017; Utterberg Modén et al., 2021).
@@
Prior work has documented instances in which teachers were able to abandon AIED deployments that failed to bring them value (e.g., Holstein et al., 2017; Utterberg Modén et al., 2021). However, even when teachers unite to refuse technology use, they are not guaranteed an escape from AI Failure Loops longer term. As discussed in Section 6.1 (judges) and Section 6.2 (doctors), higher-status workers may be able to simply ignore the outputs of an AI tool, creating natural incentives for future AI developers to account for their needs and desires in future AI creation attempts. On the other hand, lower status workers like teachers may manage to escape from an instance of a failed AI deployment, but these escapes alone are unlikely to prompt meaningful improvements in AI development practices. As discussed above, school districts and developers may still disproportionately blame teachers for low adoption and use of AIED technologies missing opportunities to identify ways to design AI technologies that genuinely support teachers.
7 Discussion: Toward Positive Futures for AI and Work

### v2

@@ 1 Introduction
Workplace AI systems are often touted by developers and deployers as tools to augment workers’ abilities and improve their work experience (Brynjolfsson et al., 2025). In practice, however, these systems have frequently fallen short of meaningfully supporting workers (Acemoglu et al., 2023). In social services, AI-based decision support tools have been introduced for high-stakes decisions, only to be withdrawn following backlash from frontline workers (Samant et al., 2021). In healthcare, clinicians ignore AI tools designed to assist in clinical decision-making (Elwyn et al., 2013; Lebovitz et al., 2022). In education, AI tutoring systems have added to teachers workloads and compromised their pedagogical autonomy—despite promises to reduce labor and enhance teaching (Holstein et al., 2017; Modén et al., 2021; Holmes, 2023). Across these sectors, AI systems fail to complement the situated, social, and tacit knowledge that underpins workers’ expertise (Lebovitz et al., 2021).
Prior research in human-computer interaction (HCI) and science and technology studies (STS) offers several explanations for these failures. Scholars have pointed to issues of poor contextual fit, where AI systems conflict with existing organizational structures and work practices (Suchman, 1987; Yang et al., 2019). Others have highlighted flawed assumptions about the nature of work, misaligned values, or inadequate governance, emphasizing how these design shortcomings intersect with broader labor and power dynamics  (e.g., Forsythe, 1993; Suchman, 1993; Star, 1999; Kawakami et al., 2024a). However, one crucial dimension remains underexplored: occupational devaluation, the systemic undermining of worker expertise, particularly within professions predominantly occupied by women and people of color (Flores-Robles and Gantman, 2024). In these professions, expertise is often not fully understood or explicitly recognized by others (Huising et al., 2025), and devalued both in status and pay even when accounting for education and experience levels (Flores-Robles and Gantman, 2024).
@@
We argue that workers engaged in devalued labor are especially vulnerable to flawed AI deployments, which often rest on reductive understandings of their work. These systems not only fail to reflect the complexity of frontline labor but can also further obscure and erode the visibility of workers’ expertise, creating a negative feedback loop that deepens occupational devaluation. Through a comparative analysis (Schwandt and Gates, 2018) of past scholarship on AI deployments across three feminized labor contexts (social work, home healthcare, K-12 teaching), we articulate why and how this dynamic emerges.
To do so, we introduce the concept of AI Failure Loops: recurring patterns of failure in the design, evaluation, and deployment of workplace AI systems, in which the devaluation of worker expertise leads to ineffective AI deployments—and those failed deployments, in turn, further erode recognition of workers’ skills. These loops are sustained across the AI development lifecycle, fueled by persistent over-estimation of what AI can do and under-recognition of the complexity and value of worker expertise.
Some of these challenges are familiar: They echo Forsythe’s observations of expert systems nearly three decades ago (Forsythe, 1993). When developing expert systems, AI developers excluded social and maintenance activities in their conceptualizations of “work,” leading to AI deployments that failed to reflect experts’ real-world work (Forsythe, 1993). Our paper examines how similar challenges have persisted with modern AI, with a focus on the role of occupational devaluation.
Today, with recent waves of both AI hype and actual increases in AI capabilities, developers are targeting a rapidly expanding range of socially complex tasks associated with feminized occupations. Drawing lessons from past AI deployments in these contexts, we can expect this will further accelerate the gross under-estimations of worker expertise and over-estimations of AI capabilities that has shaped these occupations in the past three decades. This paper responds to calls to redirect this trajectory toward pro-worker futures (e.g., Acemoglu et al., 2023), where worker expertise is celebrated within AI practice and society more broadly, and workplace AI systems are designed to truly complement worker capabilities. By centering occupational devaluation in our analysis, we extend these calls, highlighting the need for AI systems that not only avoid harm but actively contribute to the dignity and visibility of labor.
@@ 3 Devaluation of Feminized Labor
Notions of what counts as “work” and, relatedly, expertise are socially situated and have evolved across time (Suchman, 1995; Star and Strauss, 1999). These notions are used to justify the societal and economic (de)valuation of human labor. Feminized labor—which often involves care-oriented labor like teaching, home healthcare, or childcare—provides a particularly striking example. Many forms of feminized labor were historically not viewed as “work,” but as acts of love (Flores-Robles and Gantman, 2024). Even as feminized labor became formalized into paid occupations through social movements, workers struggled to gain legitimacy. Sociologists have theorized several explanations: Employers may justify low pay by conceptualizing feminized labor as “acts of love” provided out of “goodwill” (the “prisoners of love” framework (England, 2005)). The association of feminized labor with women may also lead to its devaluation precisely because women themselves are devalued in society (the devaluation theory) (England, 2005).
Even when accounting for factors like workers’ education level, occupations with a higher percentage of women pay less and have lower social status (Flores-Robles and Gantman, 2024). The most female-dominated occupations in the United States still tend to involve care-based work, including home care (85%), social work (85%), and teaching at the primary and secondary levels (77%) (Council on Social Work Education, 2023; Connacher, 2023; Taie and Lewis, 2022). Workers in feminized occupations tend to have low status in their workplace, with little say in organizational decisions impacting their day-to-day work—challenges which are exacerbated by a “glass ceiling” on career advancements (Cotter et al., 2001). These occupations often also face systemic resource shortages and high rates of worker burnout, all amid growing labor demands (England, 2005).
@@ 4.1 Team Experience and Case Selection
We examine case studies of AI deployments across social work, home healthcare, and K-12 teaching, selected because our team has direct experiences collaborating with workers in each occupation. This experience—collectively including over ten years of work observing worker practices, designing interventions with workers, and speaking with leadership and managers to facilitate these collaborations—allowed us to ground our analysis of relevant literature in our own empirical material (e.g., field notes). For example, our social work case study was informed by our prior findings from field research conducted across five U.S. public sector agencies (Kawakami et al., 2022a, 2024a). This includes experiences observing workers’ use of AI-based risk assessment tools (Kawakami et al., 2022a), speaking with agency leadership and in-house AI developers in social work agencies (Kawakami et al., 2024a, b), and designing interventions with social workers (Kawakami et al., 2022b). Our healthcare case study was informed by our experiences conducting interviews and design activities with home health aides, aide coordinators, and nurses to understand their perspectives around workplace AI (e.g., Bartle et al., 2022). Our K-12 teaching case study was informed by our prior field research studying K-12 teachers’ adoption and use of AI-based tutoring systems across hundreds of US-based schools and classrooms (e.g., Holstein et al., 2017, 2018, 2019).
@@ 4.2 Literature Review and Analysis
Our analysis is based on prior literature on AI deployments across the three case studies (social work, home healthcare, and K-12 teaching), each drawn from prior academic and grey literature (including but not limited to prior research from our team). Specifically, as we conducted a comparative case analysis (Schwandt and Gates, 2018) of AI deployments across these occupations, we began to notice common trends, which we explored further through a focused review of academic and grey literature. We gathered and analyzed literature from computing (e.g., HCI, robotics, AIED) and domain-specific (e.g., healthcare, social work, and education) research venues. We searched literature on field deployments of AI systems in workplace contexts (e.g., Jørgensen and Nissen, 2022), literature reviews on technologies deployed in specific workplaces (e.g., Kuo et al., 2022), and syntheses of trends on the impacts of AI on workers (e.g., Holmes, 2023). Our analysis was not limited to research on challenges from AI deployments. Through our initial comparative case analysis of findings across the authors, and this subsequent review of the literature, we iteratively developed the AI Failure Loops concept.
@@ 6 Case Studies
Considered part of domestic life for most families in early 19th century, while professional home care was reserved for wealthy families. By late 19th century, women began organizing efforts to send nurses to homes of poorer families (Buhler-Wilkerson, 2003). The Fair Labor Standards Act of 1938 outlaws exploitative practices and instituted an 8-hour workday but home health aides are explicitly excluded (Migda, 2015). Home health aides continue to be predominantly women and disproportionately Black or Hispanic (Connacher, 2023). Continued worker devaluation: Majority of workers lack benefits like paid time off, sick leave, or employer-sponsored health insurance (Banerjee et al., 2025); majority of workers live below poverty line (Connacher, 2023); workers often not seen as part of healthcare team, despite being integral to healthcare system (Osterman, 2018).
@@
Originally conceived as male-dominated occupation that men partook in during farming off-seasons or prior to moving into more well-established professions. Women begin to enter the profession in the late 1880s (Rury, 1991). Profession is predominantly occupied by women, except among administrators (Smith, 2022). Profession becomes increasingly formalized with state licensing and university education program requirements (tea, 2018). K-12 teachers continue to be predominantly women, except in leadership positions (e.g., principals, superintendents) which continue to be male-dominated (Kocon, 2017). Continued worker devaluation: Paid relatively little compared to other fields requiring similar levels of education (Flores-Robles and Gantman, 2024); workers report high levels of stress and burnout (Steiner and Woo, 2021).
@@ 6.1.1 Context
6.1.2 AI Failures in Social Work
@@
6.2.1 Context
Governments and private investors have thrown billions of dollars into healthcare AI innovations—and this number is only increasing each year (Esposito et al., 2016; Yang et al., 2019; Yildirim et al., 2023). Many of these innovations target home healthcare. Yet, today, there exist few, if any, technical innovations that actually support home health aides (Kuo et al., 2022)—workers, the majority of whom are women and people of color, who support elderly people and people with disabilities in everyday personal care (e.g., bathing, washing, monitoring health challenges) and household responsibilities (e.g., cooking, cleaning, shopping). Instead, existing technologies in home health have tended to automate worker skills (e.g., care robots (Esposito et al., 2016)), manage worker practices (e.g., shift matching (Solano-Kamaiko et al., 2025)), or provide support to other stakeholders like patients (e.g., interactive voice assistants (Bartle et al., 2022)). For example, to help address the immense care work shortage, several countries–across Europe, North America, and East Asia–have heavily invested in exploring the potential of care robots (Esposito et al., 2016). In fact, care robots for home healthcare applications are among the main investments in robotics applications (Esposito et al., 2016). However, a growing body of work has recognized that, if not designed carefully, these systems may disempower workers and amplify challenges they face already (Fox et al., 2019; Tseng et al., 2020; per, 2022). These tendencies to decenter home health aides in AI innovation efforts—in home health and healthcare more broadly—has occurred amid increasing acknowledgement that home health aides are the cornerstone to systems of care, that demands for their labor are increasing, and that institutional support for their wellbeing and labor has remained stagnant (Connacher, 2023).
6.2.2 AI Failures in Healthcare
@@
Even as decision-makers in healthcare (e.g., doctors, hospitals, state regulators, insurance companies) verbally acknowledge home care aides, they “typically do not think that home care aides can be real members of a care team”  (Osterman, 2019). Osterman describes this happens, in part, because home care aides’ skills are overlooked, instead misunderstood as that of “unskilled companions [and] glorified babysitters” (Osterman, 2019). Ming et. al. observes that home care aides recognize these misjudgements: “Some people see us as a maid […] We’re not just here to clean your house [..] Maid are different from an aide” (Ming et al., 2023). Ming et al. (2023) further cautions that failing to account for the invisibility of aides’ work may contribute to the deployment of AI systems that enforce compliance to impoverished notions of “care”—which, in turn, may further devalue and make invisible home health aides’ work.
@@
Researchers and policymakers have long looked for ways to sustain the growing population of elderly people—much of which is currently supported through home health aides (Bedaf et al., 2015; Hendry and Wu, 2004). While there has been limited research on technologies to support home health aides, there is a comparatively large number of papers exploring ways to automate home-based tasks needed by elderly people, including tasks home health aides usually complete (Esposito et al., 2016). Many of these technical deployments have come in the form of care robots. However, despite the massive amounts of funding dedicated towards advancing the design of care robots, actual deployments of care robots have often been deemed failures (e.g., per, 2022; Sharkey and Wood, 2014). Most of this literature has discussed failures with respect to meeting elderly people’s care needs.
While care robots may have been intended to help alleviate the increasing demand for home care workers, the few studies that account for impacts on care workers have observed impacts in the opposite direction. For example, in a scoping review of 27 papers on the deployment of care robots, researchers concluded the effectiveness of care robots is “mixed” at best. It further observed how robots can instead increase caregivers’ workload: care robots “seldom work as a shortcut to increased efficiency or time effectiveness” as workers must undergo the labor of maintaining and operating the robot (per, 2022). Another study found care workers experienced emotional demands, grappled with new ethical challenges, and felt decreased job satisfaction when care robots were deployed in their workplace (Wright, 2018, 2019).
Beyond care robots, Solano-Kamaiko et al. (2025) discuss similar challenges with other forms of home health technologies, like AI-based shift management technologies that match home health aides with patients. Home health aides reported feeling concerned about how matching algorithms fail to account for the human dimensions of their work, like personality, compassion, and empathy, that impact the quality of matching decisions (Solano-Kamaiko et al., 2025). Others described how AI-based shift management technologies may further increase the invisibility of home health aides’ work by taking away the already minimal amount of contact they have with home health agencies (Solano-Kamaiko et al., 2025).
@@
Home health aides are rarely consulted during the design process, even for technologies directly impacting them (Kuo et al., 2022). This creates a stark contrast with the development of technologies that support doctors. In this section, we briefly discuss how having organizational power as a “high status” worker can create incentives to avoid the AI Failure Loop. In doing so, we shed light onto the power and labor dynamics at play in low-status occupations, that lead to deviations in how they are impacted by AI deployments.
@@ 6.3.2 AI Failures in K-12 Teaching
There is a long history of AI development in education guided by the ultimate vision of automating away the role of human teachers (Watters, 2015; Selwyn, 2016). Over the past forty years, research and practice in AIED has largely centered students’ needs, often without explicit consideration for teachers’ needs and constraints in the classroom (Baker, 2016; Holmes, 2023; Holstein et al., 2017; Roll and Wylie, 2016). Holmes refers to these deployments as “student-focused AIED”—AI technologies aiming to automate one or more of teachers’ tasks for the stated goal of helping students learn, for example, including tutoring systems, virtual writing assistants, automatic writing evaluation, and chatbots. By contrast, Holmes finds that “teacher-focused AIED” has remained largely speculative, confined to academic research prototypes rather than commercial systems (Holmes, 2023). Beyond the research community, commercialization practices help explain the prevalence of student-focused AIED technologies at the expense of teacher-focused AIED. Companies sell AIED technologies to schools and districts, and past research has found that teachers may be given little say in whether or not to adopt a given technology (Holstein et al., 2017; Modén et al., 2021).
By reducing K-12 teachers to mere service providers responsible for operating technology and maintaining the classroom (Holmes, 2023), many commercialized AIED deployments require teachers to “accommodate” to the technology. Indeed, prior work examining AI deployments observed how teachers often struggle to make AIED technologies work in their classroom, without having to adjust their practices in unreasonable ways or take on new responsibilities along the way (Enyedy, 2014; Modén et al., 2021). Observing math teachers using an AI tutoring systems in their classrooms, Holstein et al. and Utterberg Moden et al. found that teachers were forced to grapple with numerous contradictions designed into the AIED system, eventually causing them to abandon the system. For instance, while the AI tutoring system aimed to help students work at their own pace, this ran counter to pedagogical practices around collaborative learning and discussion: Teachers had a responsibility to maintain a cohesive classroom environment, where students were roughly synchronized in the content they were learning. (Holstein et al., 2017; Modén et al., 2021).
@@
Prior work has documented instances in which teachers were able to abandon AIED deployments that failed to bring them value (e.g., Holstein et al., 2017; Modén et al., 2021). However, even when teachers unite to refuse technology use, they are not guaranteed an escape from AI Failure Loops longer term. As discussed in Section 6.1.2 (judges) and Section 6.2.2 (doctors), higher-status workers may be able to simply ignore the outputs of an AI tool, creating natural incentives for future AI developers to account for their needs and desires in future AI creation attempts. On the other hand, lower status workers like teachers may manage to escape from an instance of a failed AI deployment, but these escapes alone are unlikely to prompt meaningful improvements in AI development practices. As discussed above, school districts and developers may still disproportionately blame teachers for low adoption and use of AIED technologies missing opportunities to identify ways to design AI technologies that genuinely support teachers.
7 Discussion: Toward Pro-Worker AI